OpenStudy (anonymous):

PLEASE HELP (No calculator) Find all exact solutions to the equation sin^2x+ 2cos^2x= -2 with x between [0, 2pi]. You have an intermediate step in which you fi nd solutions to two di erent linear trigonometric equations. Identify the two equations that you used.

OpenStudy (anonymous):

trick is to rewrite \[\sin^2(x)=1-\cos^2(x)\]

OpenStudy (anonymous):

then you will have a quadratic equation in cosine

OpenStudy (anonymous):

lets just call cosine \[a\] to make the algebra easier

OpenStudy (anonymous):

\[1-a^2+2s^2=-2\] \[1+a^2=-2\] \[a^2=-3\] and either i have made a mistake, or there is no solution, or perhaps you typed it in wrong

OpenStudy (anonymous):

oh for the love of pete. the left hand side is the sum of two squares. the right hand side is negative. no solution by inspections

OpenStudy (anonymous):

sure it wasn't \[sin^2(x)+2cos^2(x)=2\]?

OpenStudy (anonymous):

Oh how confusing. These are my available answers: A) cos x = 0; tan x = 1 B) sin x = -1/2; sin x = 2 C) sin x = 0; tan x = -1 D) cos x = -1/2; cos x = 1/2 E) cos x = -1; cos x = 3

OpenStudy (anonymous):

wait wait. are you sure it is what i wrote?

OpenStudy (anonymous):

because no matter what x is \[sin^2(x)\geq0\] and likewise \[cos^2(x)\geq 0\] so there is not way to add them and get -2 !

OpenStudy (anonymous):

it definitely says -2 :( \[\sin^2 x + 2cosx =-2\] Oh gosh I am really sorry the cosine was not squared! My mistake!

OpenStudy (anonymous):

oooooooooooooooooh k

OpenStudy (anonymous):

now we proceed as before. rewrite as \[1-\cos^2(x)+2\cos(x)=-2\]

OpenStudy (anonymous):

clear?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then we get a quadratic equation in cosine \[-cos^2(x)+2cos(x)+3=0\] or \[cos^2(x)-2cos(x)-3=0\]

OpenStudy (anonymous):

like solving \[a^2-2a-3=0\] \[(a-3)(a+1)=0\] \[a = 3\] or \[a=-1\]

OpenStudy (anonymous):

now back to trig: \[cos(x)=3\] well that is impossible since the biggest cosine can be is 1

OpenStudy (anonymous):

\[cos(x)=-1\] which means \[x=\pi\]

OpenStudy (anonymous):

How do I use x= Pi to choose one of the possible answers?

OpenStudy (anonymous):

let me look at your answers A) cos x = 0; tan x = 1 B) sin x = -1/2; sin x = 2 C) sin x = 0; tan x = -1 D) cos x = -1/2; cos x = 1/2 E) cos x = -1; cos x = 3

OpenStudy (anonymous):

we started with \[sin^2(x)+2cos(x)=−2\] yes?

OpenStudy (anonymous):

number 1 is out since if cos(x)=0 there is no was sin^2(x)=-2 number 2 is out since sin(x) cannot be 2 man i am lost lost lost. maybe they just want you to write E) because that is what we got. but this doesn't make sense because there is no was cos(x) = 3

OpenStudy (anonymous):

ooh i see. it just asks for the "intermediate step" and that is what we got!

OpenStudy (anonymous):

we got E) as an "intermediate step" not clearly the final step

OpenStudy (anonymous):

the final step is to say that cos(x)=3 is impossible so cos(x)= -1 so \[x=\pi\]

OpenStudy (anonymous):

sorry but this question is confusing because you are asked to write something that is not possible. but ok it is an 'intermediate step'

OpenStudy (anonymous):

I get it! Ok so heres a douzy. The last question is written EXACTLY the same, but apparently I am supposed to solve for it a different way because the new possible answers are : A) cos x = -1; cos x = 3 B) sin x = 0; cos x = -1/2 C) cos x = 0; tan x = 1 D) sin x = -1/2; sin x = 2 E) cos x = -1/2; cos x = 1/2

OpenStudy (anonymous):

but answer A) is identical to answer E from the previous problem!

OpenStudy (anonymous):

where are you getting these?

OpenStudy (anonymous):

Uggghhhh maybe they want me to use the formula for 2cosx? its a worksheet :(

OpenStudy (anonymous):

now it is answer A) which as you can see is answer E) from before. you cannot do this a different way

OpenStudy (anonymous):

5 and 6 are the same. same problem, same answer. ignore it

OpenStudy (anonymous):

Ok, thank you!

OpenStudy (anonymous):

What is your opinion on Find all solutions of : sqrt{3}tan(x) +1 = 0 I know the solution is x = -1/sqrt{3}, and I know tangent of POSITIVE 1/sqrt{3} is pi/6.......but what does the negative do to the answer? As far as I know the solution is x= pi/6 + pik x= 5pi/6 +pik